OpenStudy (ksaimouli):

find r'(t) and |v(t)| "This is function of angle theta and t"

OpenStudy (ksaimouli):

@ikram002p @phi

OpenStudy (ikram002p):

r(t)=<v cos (theta)t+V sin(theta)t-0.5gt^2> v is diffrent in x and y ?

OpenStudy (ikram002p):

g is constant ?

OpenStudy (ksaimouli):

my bad no it is \[V0\]

OpenStudy (ksaimouli):

yes

OpenStudy (ikram002p):

type it again :3

OpenStudy (ksaimouli):

\[<v \cos \theta t,v \sin \theta t-\frac{ 1 }{ 2 } g t^2>\]

OpenStudy (ksaimouli):

The"total" question is at what angle should the projectile be fired to achieve max total distance

OpenStudy (anonymous):

this is a position vector with respect to t and theta?

OpenStudy (phi):

my first thought is maximize r(t) dot r(t) (square of the distance)

OpenStudy (ikram002p):

so is it r(t) ?

OpenStudy (anonymous):

you can try using partial derivatives with respect to maximize distance as a function of t.

OpenStudy (anonymous):

with respect to theta*

OpenStudy (ksaimouli):

my professor gave the answer on how to solve. She said "v(t) = r ' (t) = ..... Find | v(t) |. This is function of angle @ and t. Now it hits the ground gives t = (2 v/ g) sin @. Actually this gives limits of t for the entire distance travelled from t = 0 to (2v/g) sin @, when it completes its path. So the distance travelled by the projectile say L(@) , is given by L(@) = Integral of | v(t) | dt within limits 0 to (2v/g) sin @. You may use the tables at the back of the book for integration purpose. IT will come out to bev^2/ g sin @ + v^2/ 2g cos ^2 @ ln [ (1+ sin @)/ (1 - sin @) ] to maximize L9@) for 2 between 0 and 90 degrees, L '(@) = 0, gives @ nearly 0.9855. Compare the L value at this Angle and end points 0 and 90 degress. Max value of L (1.20 v^2/g) is at this angle ).9855 which is nearly 56 degrees"

OpenStudy (ksaimouli):

I don't understand why we r taking derivative and then taking integral also is this partial derivative?

OpenStudy (anonymous):

because you want to maximize the function for t while keeping theta variable...

OpenStudy (anonymous):

but this might be a little rough...actually I think I was confused on what your vector represented. Is it the position vector?

OpenStudy (ksaimouli):

yes

OpenStudy (anonymous):

do you want max arclength or distance in x direction.

OpenStudy (phi):

I assume only theta is an argument to cos or sin? not theta*t in other words, t is multiply cos(theta) ?

OpenStudy (ksaimouli):

yes

OpenStudy (ksaimouli):

\[r(t)=(v \cos \theta t)i+[v \sin \theta t-\frac{ 1 }{ 2 }g t^2]j\]

OpenStudy (ksaimouli):

from the knowledge of physics g= gravity and t = time

OpenStudy (phi):

it's less confusing if you write it as \[ r(t)=(v \ t \cos \theta )i+[v \ t \sin \theta -\frac{ 1 }{ 2 }g t^2]j \]

OpenStudy (ikram002p):

well since its r(t) then it wud be one variable fun :-\ idk theta situation sounds weird xD

OpenStudy (phi):

what do you mean by maximum total distance ? is this arc length or horizontal distance ?

OpenStudy (anonymous):

the reason you need to take the derivative of the integral is the total distance is modeled by the the integral of the speed. Speed is derivative of the norm. so you would use the arc length formula. (also I am confused about the variables as well)

OpenStudy (ksaimouli):

so \[v(t)= <v \cos \theta, v \sin \theta - g t>\]

OpenStudy (anonymous):

this is the velocity vector.

OpenStudy (ksaimouli):

ya

OpenStudy (anonymous):

first clarify if you mean arc length vs horizontal distance please

OpenStudy (ksaimouli):

I dont know what professor meant by total distance

OpenStudy (ksaimouli):

Integral of | v(t) | dt within limits 0 to (2v/g) sin @.

OpenStudy (anonymous):

thats defintely arclength formula....so you need to first find the derivative of the position vector and get the norm of that. This is the speed. next you integrate that with respect to t to get arclength formula.

OpenStudy (anonymous):

I think from there you can maximize the function with respect to theta...but I am not sure.

OpenStudy (ksaimouli):

\[\int\limits_{0}^{\frac{ 2v \sin \theta }{ g }}\sqrt{(vsin \theta)^2+(vsin \theta-g t)^2}\]

OpenStudy (ksaimouli):

right?

OpenStudy (anonymous):

instead of sin, the first one is cos

OpenStudy (ksaimouli):

lol yes thx, \[v^2/ g \sin \theta + v^2/ 2g \cos ^2 \theta \ln [ (1+ \sin \theta)/ (1 - \sin \theta) ]\]

OpenStudy (ksaimouli):

i need to take derivative of this and find critical points?

OpenStudy (ksaimouli):

What is whole point of this I mean why are we useing arc length

OpenStudy (anonymous):

because you prof told you to take integral of norm of derivative, which is the formula for arc length.

OpenStudy (ksaimouli):

so i got 0.98551474 how to get the angle?

OpenStudy (phi):

you found theta in radians. Convert to degrees (multiply by 180/pi) sylbot outlined what you are doing: your position vector \[ r(t)=(v\ t\cos \theta )i+[v \ t\sin \theta-\frac{ 1 }{ 2 }g t^2]j \] traces out an arc. the velocity at any time t is r'(t) \[ r'(t)= <- v \sin \theta , \ v \cos \theta - g t > \] which is a vector with magnitude and direction The speed is the magnitude of the velocity vector If the speed were constant, the distance traveled would be speed * time but when the speed s(t) varies we integrate \[ \int s(t) dt \]

OpenStudy (phi):

to find the bounds for the integral, we assume the arc begins and ends at height zero, or in other words, with the y-component of the position vector= 0 solving \[ v \ t\sin \theta-\frac{ 1 }{ 2 }g t^2 = 0 \] for t we get t=0 and t= (2v/g) sin theta

OpenStudy (phi):

after integrating you have found the length of the arc as a function of the launch angle theta. Take the derivative with respect to theta to find the critical points (max in this case) of the length of the arc. you find theta= 0.98551474 radians or 56.47 degrees

OpenStudy (phi):

if you use theta= 0.98551474 in the formula for the length of the arc \[ \frac{v^2}{ g} \sin \theta + \frac{v^2}{2 g} \cos ^2 \theta \ \ln [ (1+ \sin \theta)/ (1 - \sin \theta) ] \] and simplify , you get \[ 1.1996 \frac{v^2}{ g} \] compare that to using theta= 0 or theta=pi/2